The length of a rectangle is 2 centimeters less than three times its width.If the area is 40 square centimeters,find the length and width.
Answer by checkley71(8403)

L=3W-2
L*W=40
(3W-2)W=40
3W^2-2W-40=0
(3W+10)(W-4)=0
3W+10=0
3W=10
W=-10/3 NOT A SOLUTION
W-4=0
W=4
L=3*4-2
L=12-2
L=10
PROOF
10*4=40
40=40